Rice planting machine Swami Kalekar:
Status: Rice sapling for trial have arrived. The major issue of the simultaneous movement of delivery hand and backward movement of the shaft is yet unsolved. A round of discussion with Mr. Vijay Kumar, Mr. Ashok Kalbag and Sandeep Shinde resulted in to new set of ideas. Not yet finalized on a single one.Magnet driven device was indicated to be problematic during these discussion. Action Item: Not clear.
Basundi Maker: Supriya Kadam
Status:The basudi maker is expected to go for 1st real life trail on 9 Jan 2017. The provisos trail done in the 1st week of Jan resulted in to loss of 350 ml residual water from 3 lit loaded in 2 .5 hrs. During his trail AC current was used to run the motor. Action Items:Next trail is expected to be with DC current generated by photo voltaic cell. The other problem faced was softenting of acrylic sheet used for motor box.
Non-electric Oxygen Enrichment Unit : Dipali Kambale/ Subir:
Status No change in status in this week. Trail on the first version is complete. Pressure 5 kg and flow 40 m/ lit achieved. Issue : (1)MOC of the wheel continued to be challenge rubber bushing were fixed on the nylon wheel with help of screws. These came out due to stress. (2) The frame was modified to reduced vibration. Very tight v belt resulted in high friction. Loosening of the belt resulted in brushing of belt with metal frame. Needs 0.5 mm elevation for compressor as well as bearing .
Polyhouse controller / Bio-Dome Suhas Labde:
Status : Calibration for air humidity meter done. 200 lit water used per day due to siphoning of water even thought drip motor was off. Higher temperature noted and hence exhaust fan was activated under two conditions (1)temperature higher than 30 and (2) air humidity more than 60%. Conditions seems to be operating but temp of 34oC was note in the afternoon on 7 Jan 17. One degree temperature rise was observed in 4 min when exhaust was put off . Humidity was observed to be 28% where as out side humidity is 16 according to google. Exhaust fan with higher capacity or cooling pad necessary to control temp below 35 if necessary.
Hackberry hand Prachi Samarth
Status : Hackberry hand was tested in the first week of Jan 2017. It is working and need some more improvement in the next version.
Ambient temperature dryer Shubham Shembade
Sataus : Dome sealed with cement and ready to use. Turbo ventilator purchased installation is pending. Action Items: Need to obtain temperature and flow profile with and with out ventilator and with and with out trays. Two set of trays control and experimental at north and south side of dome with and with out silica to generate drying time and humidity related data
Shredder New project assigned.: Vishwas Shinde
Design ready but not shared with Subir or team. Project is facing issue due to non functionality of plasma cutter.
Bio-drum Ameya Kulkani/ Arun Dixit:
Status: the 2nd version not ready. Necessary hardware is on site. Issue is alignment of sprocket wheel. Action Item: Aligning the wheel to get motorized drum oerational by 14 Jan 17.
Signal buster Ravi
- 4 Orders with advance booked for Signal Buster.
- Expected to get material by 10th Jan 17 and expected to get sale the product with in a weeks time.
Mr. |Ankit Panchabhai
Mr. Ankit Panchabhai has done initial market survey of existing desert cooler inthe market with respect to the fan and pump as well as size.
Ganesh Nikhare: Left DIC review will be discontinued form next cycle.